OpenStudy (anonymous):

what is the alternating sequence of e^-n what is the alternating sequence of e^-n @Mathematics

OpenStudy (anonymous):

power series expansion o f \[e^x\] is \[\sum_{k=0}^{\infty}\frac{x^k}{k!}\] replace x by -x to get \[\sum_{k=0}^{\infty}\frac{(-x)^k}{k!}\] or if you prefer \[\sum_{k=0}^{\infty}\frac{(-1)^kx^k}{k!}\]

OpenStudy (anonymous):

does that mean its geometric?

OpenStudy (anonymous):

no it is not geometric. geometric means there is a common ratio, as in \[\sum ar^n\]

OpenStudy (anonymous):

so the alternating sequence DNE?

OpenStudy (anonymous):

?